OpenStudy (anonymous):

Find the average value of f(x) = sin(x) over the interval [0, π].

jimthompson5910 (jim_thompson5910):

Average value (A) of a function f(x) over the interval from x = a to x = b \[\Large A = \frac{1}{b-a} \int_{a}^{b} f(x) dx\]

OpenStudy (anonymous):

pi/2?

jimthompson5910 (jim_thompson5910):

In your case, you have a = 0 b = pi f(x) = sin(x) \[\Large A = \frac{1}{b-a} \int_{a}^{b} f(x) dx\] \[\Large A = \frac{1}{\pi-0} \int_{0}^{\pi} \sin(x) dx\] \[\Large A = \frac{1}{\pi} \int_{0}^{\pi} \sin(x) dx\] \[\Large A = ???\]

jimthompson5910 (jim_thompson5910):

what's the integral of sin(x) ?

OpenStudy (anonymous):

-cos

jimthompson5910 (jim_thompson5910):

yes -cos(x) + C we can drop the +C because we have a definite integral

jimthompson5910 (jim_thompson5910):

evaluate -cos(x) when x = pi evaluate -cos(x) when x = 0 subtract the results to get ???

OpenStudy (anonymous):

2

OpenStudy (anonymous):

I got pi over 2 but someone told me just 2

OpenStudy (anonymous):

or sorry 2 over pi

jimthompson5910 (jim_thompson5910):

yeah the integral portion should evaluate to 2

jimthompson5910 (jim_thompson5910):

which is why A = 2/pi

OpenStudy (anonymous):

okay thank you very much for clearing that up

jimthompson5910 (jim_thompson5910):

np